A car traveled from A to B at 50 kph, from B to C at 60 kph, and returned(C to B to A) at 80 kph. What was the average speed (1) on the outbound trip and (2) on the round trip if the distances from A to B and B to C were, respectively, a. 100 km and 120 km?

Found 2 solutions by math_tutor2020, josgarithmetic:
Answer by math_tutor2020(3816)


Answers:
(1) 55 kph
(2) 65.185 kph (approximate)

======================================================================

Explanation:

Part (1)

Here's one way we can draw out the diagram to represent what's going on:


For the portion from A to B, we have
distance = 100 km
rate = 50 kph

Let's determine how long it takes to go from A to B
distance = rate*time
time = distance/rate
time = (100 km)/(50 kph)
time = (100/50) hrs
time = 2 hrs

Then from B to C
time = distance/rate
time = (120 km)/(60 kph)
time = (120/60) hrs
time = 2 hrs

The total outbound trip (A to B to C) takes 2+2 = 4 hours.
The total distance is 100 km + 120 km = 220 km.

Then,
distance = rate*time
rate = distance/time
rate = (220 km)/(4 hrs)
rate = (220/4) kph
rate = 55 kph

----------------------------------------------

Part (2)

When returning home (C to B to A), we're now going 80 kph and we travel 220 km.

time = distance/rate
time = (220 km)/(80 kph)
time = (220/80) hrs
time = 2.75 hrs

The outbound time previously calculated was 4 hours.
The return time is 2.75 hrs
The total round trip time is 4+2.75 = 6.75 hrs

The total round trip distance is 2*220 = 440 km

rate = distance/time
rate = (440 km)/(6.75 hrs)
rate = (440/6.75) kph
rate = 65.185185185185 kph
That value is approximate. The "185"s repeat forever.
Round that however your teacher instructs.

If we were to round to 3 decimal places, then the round trip average rate is roughly 65.185 kph


Answer by josgarithmetic(39615)

ROUTE      SPEED(kilom/hr)      TIME           DISTANCE
A to B       50                 100/50           100
B to C       60                 120/60           120
C to A       80                 220/80           220

That make sense?
Decide what to do wit it?